# Eating Speed Measurement Using Wrist-Worn IMU Sensors in Free-Living Environments

### v1

@@ Abstract
Eating speed is an important indicator that has been widely scrutinized in nutritional studies. The relationship between eating speed and several intake-related problems such as obesity, diabetes, and oral health has received increased attention from researchers. However, existing studies mainly use self-reported questionnaires to obtain participants’ eating speed, where they choose options from slow, medium, and fast. Such a non-quantitative method is highly subjective and coarse in individual level. In this study, we propose a novel approach to measure eating speed in free-living environments automatically and objectively using wrist-worn inertial measurement unit (IMU) sensors. Specifically, a temporal convolutional network combined with a multi-head attention module (TCN-MHA) is developed to detect bites (including eating and drinking gestures) from free-living IMU data. The predicted bite sequences are then clustered to eating episodes. Eating speed is calculated by using the time taken to finish the eating episode to divide the number of bites. To validate the proposed approach on eating speed measurement, a 7-fold cross validation is applied to the self-collected fine-annotated full-day-I (FD-I) dataset, and a hold-out experiment is conducted on the full-day-II (FD-II) dataset. The two datasets are collected from 61 participants in free-living environments with a total duration of 513 h, which are publicly available. Experimental results shows that the proposed approach achieves a mean absolute percentage error (MAPE) of 0.110 and 0.146 in the FD-I and FD-II datasets, respectively, showcasing the feasibility of automated eating speed measurement. To the best of our knowledge, this is the first study investigating automated eating speed measurement.
Index Terms:
@@ 1 Introduction
Eating speed is considered as an important factor associated with body mass index (BMI), obesity and diabetes, which has been widely investigated [1, 2, 3]. Participants with faster eating speeds are considered more likely to have higher BMI, a higher risk of obesity, diabetes, and cardiovascular disease. Furthermore, deviations in eating speed are also correlated with eating disorders [4]. Currently the most popular method for investigating eating speed is through self-reported questionnaires [2, 3]. In the work of Kudo et al.[3], participants were asked to choose their own eating speed from options such as slow, medium, and fast, by responding to questions like “How fast do you eat compared to others around same ages? (Faster, Normal, Slower).” The questionnaire based eating speed estimation is highly subjective, and there is no standard reference to define an appropriate objective eating speed. While self-reported eating speed may be sufficient at a group level, especially with a large population, it is an unreliable approach to assess an individual’s eating speed, particularly in the context of precision healthcare[5]. There is a call for an automated and objective approach to measure eating speed.
@@
Recently, automated food intake monitoring has drawn lots of attention, plenty of approaches haven been proposed to detect bites [6, 7, 8, 9] during meal sessions, detect eating episodes in free-living scenarios[10, 11], and estimate calorie intake using various sensors (e.g., inertial, camera, microphone, proximity) and machine learning techniques. However, to date, there has been no research that focus on automated eating speed detection in free-living environments.
@@
In this study, we use the term bite to refer to eating gestures and drinking gestures. The definitions of eating gestures and drinking gestures are consistent with the work in [12]. Specifically, they are defined as the action of raising the hand to the mouth with cutlery or a water container until the hand is moved away from the mouth. The definition of objective eating speed is the number of bites divided by the time taken to finish the eating episode (bites/min) [1, 13]. Based on this definition, a straightforward approach for eating speed estimation is to combine the bite detection (to count the number of bites) and eating episode localization (to obtain the time duration of the consumed eating episode). However, the reason that hinder automated eating speed estimation is two-fold: Firstly, existing bite detection approaches only focus on in-meal scenarios, it is challenging to detect bites in free-living environments. Secondly, current eating episode localization approaches cannot precisely segment the boundary of detected eating episode, whereas eating speed detection requires accurate boundaries detection for each eating episode.
In this research, we extend the bite detection from in-meal scenarios to free-living environments, and further cluster the detected bite sequences into eating episodes, to facilitate the automated eating speed estimation. The main contributions of this research can be summarized as follows:
@@
  • A complete framework for eating speed measurement in free-living environments is proposed: A sequence-to-sequence (seq2seq) temporal convolutional network combined with a multi-head attention (TCN-MHA) model is designed to process inertial measurement unit (IMU) data for detecting food intake gestures and segmenting the time interval of intake gestures in free-living environments. The obtained bite sequences are clustered into eating episodes to calculate eating speed. To our best knowledge, this is the first work to automatically estimate eating speed in free-living environments.
@@
  • An intensive comparison between our approach and existing works has been implemented. Five existing deep learning models are implemented for comparison. Additionally, the proposed eating speed measurement method is validated in two studies: 7-fold cross validation on the well-annotated FD-I dataset, and a hold-out validation on the coarsely annotated FD-II dataset.
@@
  • We make two datasets collected in this study publicly available [^1] We plan to add the link of this dataset in the revision stage.. Specifically, the FD-I dataset contains IMU data collected from 34 participants in free-living environments with fine annotation. This is the first full-day IMU dataset that contains eating and drinking gesture annotations not only during meal sessions, but also outside of meal sessions. The FD-II dataset serves as a hold-out dataset, containing IMU data from 27 participants in free-living environments.
@@ 2 Related work
Automated eating speed detection relies on two essential tasks: bite detection to count the number of bites and eating episode detection to predict the duration of each meal. In this section, we firstly introduce existing approaches for in-meal bite detection; then, we discuss eating episode detection approaches. Thirdly, we present a few studies measuring eating speed objectively (but not automated). Finally, deep learning models for time-series signal processing are introduced.
@@ 2.1 Bite Detection
Bite detection has been widely investigated using various sensors. Cameras have been used to detect bites [14, 15] and food types [7] by analyzing video signals. Acoustic sensors can be used for this task through processing chewing sounds [8]. Mertes et al. [16] developed a strain gauge-based smart plate to detect bites based on the weight change of food. In our recent work [12], a novel FMCW radar-based system was proposed for in-meal bite detection. This system was validated using our public Eat-Radar dataset, which includes 70 meals from 70 participants. The photoplethysmography (PPG) sensor [17] and electromyography (EMG) sensor [18] have also been explored for bite detection. Apart from these sensors, to date, the wrist-worn IMU sensor is a popular choice for bite detection because of its least burdensome and most acceptable. Dong et al. [19] developed a rule-based approach to detect bite using the rotation velocity of wrist. Shen et al. [20] further evaluated Dong’s approach on Clemson dataset containing 488 eating episodes. Kyritsis et al. [6] proposed an end-to-end based approach using a convolutional neural network and long-short-term-memory network hybrid model (CNN-LSTM) model to detect bite automatically on FIC dataset. Rouast et al. [21] further developed single-stage ResNet based CNN-LSTM architecture for bite detection on the OREBA dataset, which contains 100 meals. Wei et al.[22] developed an energy-efficient approach, specifically, an optimized multicenter classifier (O-MCC) and an Android application, to detect intake gestures with low inference time.
The aforementioned bite detection approaches have shown promising performances. However, it should be noted that these approaches only focus on bite detection in-meal sessions (10-20 min). Although such in-meal detection is the key and fundamental element in food intake monitoring, a more challenging but meaningful scenario, bite detection in free-living environments ($\geq{6}$ h), has yet to be broadly investigated. There are several obstacles impeding the detection in free-living environments. Firstly, it is troublesome to obtain bite-level annotation information outside meal sessions. Secondly, in the scale of full-day duration, bites are extremely sparse, leading a highly imbalanced dataset compared to in-meal datasets, making bite detection more challenging.
Figure 1: Examples of wrist-worn IMU data collection. Two IMU sensors were mounted on both hands. Participants completed their daily activities without restriction. This figure only shows the food intake related scenes, their other daily activities, such as studying, walking, talking, running, cooking, were recorded by IMU sensor as well.
@@ 2.2 Eating Episodes Detection
Eating episodes detection is another popular research topic in food intake monitoring. Such a system mainly focus on the detection of eating episodes under free-living environments. A standard process pipeline involves cutting the full-day data into minute-level segments, and machine learning is used to predict if each segment belongs to eating episode or not. Sharma et al. [10] collected a data set contains 354 days of 6-axis IMU data from 351 subjects. The IMU sensor were mounted on the dominant wrist. Doulah et al. [11] developed the AIM-2 system, a pair of eyeglasses mounted with a camera, and a 3-axis accelerometer to detect eating episodes. The AIM-2 was validated on a dataset collected from 30 volunteers.
Unlike the above methods that directly predict eating episodes, another routine is first to detect basic elements of eating episode such as chewing, swallowing, and hand-to-mouth events, and then combine them together as an eating episode using various merging techniques. Bedri et al. [23] proposed the FitByte eyeglass-based system to detect eating/drinking events and eating episodes. They detected the episodes by merging any detected intakes that are within 5 minutes from each other. Zhang et al. [24] developed Necksense system to detect chewing sequences, then clustered the detected sequences into eating episodes using a density-based spatial clustering of applications with noise (DBSCAN) algorithm [25]. Kyritsis et al. [6] first detected bites on FreeFIC dataset, then applied a Gaussian filter on the bite sequence to generate meal regions. It should be noted that the model used to detect bites is trained on in-meal FIC dataset, there is no bite-level label in FreeFIC dataset, so the bite detection performance is unclear on FreeFIC dataset.
@@ 2.3 Eating Speed Estimation
Two studies have been found to objectively (but not automated) measure eating speed. Woodward et al. [5] provided a 550 g meal to each participant and used a stopwatch to record the meal time. Their experiment was conducted in lab environments. Alshurafa et al. [13] measured eating speed by using wearable fish-eye camera in free-living environments. A trained annotator manually counted the number of intake gestures, selected the boundaries of meal sessions by viewing the video. The eating speed was then calculated by using meal duration to divide the number of bites. The aforementioned approaches require manual processing, which are time-consuming and labor-intensive.
@@ 2.4 Temporal Sequence Models
The commonly used model in intake gesture detection is CNN with recurrent neural networks (CNN-RNNs). In this architecture, CNN is typically used to extract features from time-series data, and the feature sequences are then fed into RNNs to process temporal dependencies. Although RNNs can effectively process time-series data, they struggle to memorize long-term interdependencies due to the gradient vanishing problem. Two solutions have been proposed in the literature: the temporal convolutional network (TCN) and self-attention.
@@ 2.4.2 Self-Attention
The self-attention module was proposed by Vaswani et al. [29] to compose the transformer architecture in natural language processing (NLP) domain, showcasing its superior ability. Motivated by its success, several approaches have been proposed by integrating the self-attention mechanism into CNN and RNN architectures to further improve the capability of time-series signal modeling [30, 31].
@@ 3.1 Sensors
As this experiment aims to record data in free-living environments, food intake events can happen with both hands, thus, two shimmer3 IMU wristbands [^2] https://shimmersensing.com/product/shimmer3-imu-unit/ were mounted on both hands of participants. The battery duration of shimmer3 IMU is 24 h, which satisfies the requirement of this experiment. The sampling frequency was set to 64 Hz. The 3-axis accelerometer and 3-axis gyroscope units were activated to generate 6 channels IMU data. The data were stored in SD card of shimmer, and can be downloaded into laptop via Consensys software [^3] https://shimmersensing.com/product/consensyspro-software/ for further data processing. A camera was used to record the experiment for annotation. The sensor deployment and data collection example are shown in Fig. 1.
@@ 3.2 Full-Day Data Collection
This research was approved by the ethical committee of KU Leuven with project number: G-2021-4025-R4. Informed consent was obtained from all participants. Two datasets were collected for this study, namely the well-annotated full-day-I (FD-I) dataset, and hold-out full-day-II (FD-II) dataset. There is no participants overlap among the two datasets.
@@ 3.2.1 FD-I Dataset
The FD-I dataset contains 34 days of IMU data from 34 participants (6 of them are from our previous study on drinking activity detection [32]). On the data collection day, our research assistants met the participant, instructed the participant to wear IMU wristbands. Participants were free to engage in their normal daily activities. The daily activity of each participant was recorded by a camera for annotation. Experiment locations contained participant’s home (apartments, student residence), restaurants, library, university learning center, and campus rest areas. Our research assistants were responsible for the recording when participants change their locations to ensure that all eating and drinking gestures were captured. There were no limitations on the participants’ activity during data collection. At least one meal was collected from each participant, and the minimum data collection duration was 6 h. Both eating alone and social eating scenarios were included in the dataset. The participants received restaurant voucher (20 euro) as experiment compensation after the data collection. A total of 251 h two-hand IMU data were collected, which contains 74 eating episodes, with 4,568 eating and 1,100 drinking gestures. The dataset contains four eating styles including forks & knives, chopsticks, spoon, and hands.

### v2

# Eating Speed Measurement Using Wrist-Worn IMU Sensors Towards Free-Living Environments

@@ Abstract
Eating speed is an important indicator that has been widely investigated in nutritional studies. The relationship between eating speed and several intake-related problems such as obesity, diabetes, and oral health has received increased attention from researchers. However, existing studies mainly use self-reported questionnaires to obtain participants’ eating speed, where they choose options from slow, medium, and fast. Such a non-quantitative method is highly subjective and coarse at the individual level. This study integrates two classical tasks in automated food intake monitoring domain: bite detection and eating episode detection, to advance eating speed measurement in near-free-living environments automatically and objectively. Specifically, a temporal convolutional network combined with a multi-head attention module (TCN-MHA) is developed to detect bites (including eating and drinking gestures) from IMU data. The predicted bite sequences are then clustered into eating episodes. Eating speed is calculated by using the time taken to finish the eating episode to divide the number of bites. To validate the proposed approach on eating speed measurement, a 7-fold cross validation is applied to the self-collected fine-annotated full-day-I (FD-I) dataset, and a holdout experiment is conducted on the full-day-II (FD-II) dataset. The two datasets are collected from 61 participants with a total duration of 513 h, which are publicly available. Experimental results show that the proposed approach achieves a mean absolute percentage error (MAPE) of 0.110 and 0.146 in the FD-I and FD-II datasets, respectively, showcasing the feasibility of automated eating speed measurement in near-free-living environments.
{IEEEkeywords}
@@ 1 Introduction
Eating speed is considered as an important factor associated with body mass index (BMI), obesity and diabetes, which has been widely investigated [1, 2, 3]. Participants with faster eating speeds are considered more likely to have higher BMI, a higher risk of obesity, diabetes, and cardiovascular disease. Furthermore, deviations in eating speed are also correlated with eating disorders [4]. Currently the most popular method for investigating eating speed is through self-reported questionnaires [2, 3]. In the work of Kudo et al. [3], participants were asked to answer questions like “How fast do you eat compared to others around same ages? (Faster, Normal, Slower).” The questionnaire based estimation is highly subjective, and there is no standard reference to define an appropriate objective eating speed. While self-reported eating speed may be sufficient at a large group level, it is an unreliable approach to assess an individual’s eating speed, particularly in the context of precision healthcare[5]. There is a call for an automated and objective approach to measure eating speed.
@@
Recently, automated food intake monitoring has drawn lots of attention, plenty of approaches haven been proposed to detect bites [6, 7, 8, 9] during meal sessions, detect eating episodes in free-living scenarios[10, 11] using various sensors (e.g., inertial, camera, microphone, proximity). However, to date, there has been no research that focus on automated eating speed detection in (near-)free-living environments.
@@
In this study, we use the term bite to refer to eating and drinking gestures. The definitions of eating and drinking gestures are consistent with the work in [12]. Specifically, they are defined as the action of raising the hand to the mouth with cutlery until the hand is moved away from the mouth. The definition of objective eating speed is the number of bites divided by the time taken to finish the eating episode (bites/min), which has been utilized in [1, 5, 13]. Based on this definition, a straightforward approach for automated eating speed estimation in full-day scenario is to combine the bite detection (to count the number of bites) and eating episode localization (to obtain the time duration of the consumed eating episode). However, the reason that hinder automated eating speed estimation is two-fold: Firstly, existing bite detection approaches only focus on in-meal scenarios, it is challenging to detect bites in (near-)free-living environments. Secondly, current eating episode localization approaches cannot precisely segment the boundary of detected eating episode, whereas eating speed detection requires accurate boundary segmentation.
In this study, we combine two classical tasks in food intake monitoring, i.e., bite detection and episode detection, to facilitate the automated eating speed estimation. The main contributions of this research can be summarized as follows:
@@
  • A complete framework for eating speed measurement in near-free-living environments is proposed: A sequence-to-sequence (seq2seq) temporal convolutional network combined with a multi-head attention (TCN-MHA) model is designed to process inertial measurement unit (IMU) data for detecting food intake gestures and segmenting the time interval of intake gestures. The obtained bite sequences are clustered into eating episodes to calculate eating speed.
@@
  • An intensive comparison between our approach and existing works has been implemented. The proposed eating speed measurement method is validated in two studies: 7-fold cross validation on the well-annotated full-day-I (FD-I) dataset, and a holdout validation on the coarsely annotated full-day-II (FD-II) dataset. Additionally, the public OREBA and Clemson are utilized to further analyze the performance of proposed approach. To our best knowledge, this is the first work to automatically estimate eating speed in near-free-living environments.
@@
  • We make two datasets collected in this study publicly available [^1] https://rdr.kuleuven.be/dataset.xhtml?persistentId=doi:10.48804/CN8VBB. Specifically, the FD-I dataset contains IMU data collected from 34 participants in near-free-living environments with fine annotation. This is the first full-day IMU dataset that contains eating and drinking gesture annotations not only during meal sessions, but also outside of meal sessions. The FD-II dataset serves as a holdout dataset, containing IMU data from 27 participants in free-living environments.
@@ 2 Related work
Eating speed detection relies on two essential tasks: bite detection to count the number of bites and eating episode detection to predict the duration. In this section, we firstly introduce existing approaches for in-meal bite detection; then, we discuss eating episode detection approaches. Thirdly, we present a few studies measuring eating speed objectively. Finally, deep learning for time-series signal are introduced.
@@ 2.1 Bite Detection
Bite detection has been widely investigated using various sensors. Cameras have been used to detect bites [14, 15] and food types [7]. Acoustic sensors can be used for chewing sounds detection[8]. Mertes et al. [16] developed a strain gauge-based smart plate to detect bites based on the weight change of food. In our recent work [12], a novel radar-based system was validated using our public Eat-Radar dataset. The photoplethysmography (PPG) sensor [27] and electromyography (EMG) sensor [18] have also been explored for bite detection. Apart from these sensors, to date, the wrist-worn IMU sensor is a popular choice for bite detection because of its least burdensome and most acceptable. Dong et al. [19] developed a rule-based approach to detect bite using the rotation velocity of wrist. Shen et al. [20] further evaluated Dong’s approach on Clemson dataset. Kyritsis et al. [6] proposed an end-to-end based approach using a convolutional neural network and long-short-term-memory network (CNN-LSTM) model to detect bite on FIC dataset. Rouast et al. [21] further developed single-stage ResNet based CNN-LSTM architecture for bite detection on the OREBA dataset. Wei et al. [22] developed an energy-efficient approach integrating an optimized multicenter classifier (O-MCC) to detect intake gestures with low inference time.
The aforementioned approaches have shown promising performances. However, it should be noted that these approaches only focus on bite detection within meal sessions (10-20 min). A more challenging scenario, bite detection in (near-)free-living environments ($\geq{6}$ h), has yet to be broadly investigated. There are several obstacles impeding the detection. Firstly, it is troublesome to obtain bite-level label outside meal sessions. Secondly, in the scale of full-day duration, bites are extremely sparse, leading a highly imbalanced dataset compared to in-meal datasets, making bite detection more challenging.
Figure 1: Examples of wrist-worn IMU data collection. Two IMU sensors were mounted on both hands. Participants completed their daily activities without restriction. This figure only shows the food intake related scenes, their other daily activities, such as studying, walking, talking, running, cooking, were recorded by IMU sensors as well.
Figure 2: (a) the data collection protocol; (b) the annotation method, the video label is for state 1, and the all ’0’ label is for state 2.
@@ 2.2 Eating Episodes Detection
Eating episodes detection is another popular research topic in food intake monitoring. Such a system mainly focus on the detection of eating episodes under free-living environments. A standard process pipeline involves cutting the full-day data into minute-level segments, and machine learning is used to predict if each segment belongs to eating episode or not. Sharma et al. [10] collected a dataset including 354 days of wrist-worn IMU data from 351 subjects. Doulah et al. [11] developed the AIM-2 system, a pair of eyeglasses mounted with a camera, and a 3-axis accelerometer to detect eating episodes.
Unlike the above methods that directly predict eating episodes, another routine is first to detect basic elements of eating episode such as chewing, swallowing, and hand-to-mouth events, and then combine them together as an eating episode using various merging techniques. Bedri et al. [23] proposed the FitByte eyeglass-based system to detect intake events and eating episodes. They detected the episodes by merging any detected intakes that are within 5 minutes from each other. Zhang et al. [24] developed Necksense system to detect chewing sequences and cluster detected sequences into eating episodes using a density-based spatial clustering of applications with noise (DBSCAN) algorithm [25]. Kyritsis et al. [6] first detected bites on FreeFIC dataset, then applied a Gaussian filter on the bite sequence to generate meal regions. Noted that the model used to detect bites is trained on in-meal FIC dataset, there is no bite-level label in FreeFIC dataset, so the bite detection performance is unclear on FreeFIC dataset.
@@ 2.3 Eating Speed Estimation
Several studies have been found to objectively (but not automated) measure eating speed. Woodward et al. [5] provided a 550 g meal to each participant and used a stopwatch to record the meal time in lab environments. Scisco et al. [26] explored the utility of the Bite Counter device during meal sessions, and further utilized the device in free-living environments to estimate eating speed [27]. Notably, in [27], users were required to press a button to start and stop the counter manually, and the accuracy of bite detection was unclear due to the absence of ground truth bite information. Alshurafa et al. [13] measured eating speed by using wearable fish-eye camera in free-living environments. A trained annotator manually counted the number of intake gestures, selected the boundaries of meal sessions by viewing the video. The eating speed was then calculated by using meal duration to divide the number of bites.
@@ 2.4 Temporal Sequence Models
The commonly used models in intake gesture detection are CNN with recurrent neural networks (CNN-RNNs). The CNN is typically used to extract features from time-series data, and the feature sequences are then fed into RNNs to process temporal dependencies. Although RNNs can effectively process time-series data, they struggle to memorize long-term interdependencies due to the gradient vanishing problem. Two solutions have been proposed in the literature: the temporal convolutional network (TCN) and self-attention.
@@ 2.4.2 Self-Attention
The self-attention module was proposed by Vaswani et al. [31] to compose the transformer architecture in natural language processing (NLP) domain. Motivated by its success, several approaches have been proposed by integrating the module into CNN and RNN architectures to further improve the capability of time-series signal modeling [32, 33].
@@ 3.1 Sensors
As this experiment aims to record data in (near-)free-living environments, food intake events can happen with both hands, thus, two shimmer3 IMU wristbands [^2] https://shimmersensing.com/product/shimmer3-imu-unit/ were mounted on both hands of participants. The battery duration of shimmer3 IMU is 24 h, which satisfies the requirement of this experiment. The sampling frequency was set to 64 Hz. The 3-axis accelerometer and 3-axis gyroscope units were activated to generate 6 channels IMU data. The data were stored in SD card of shimmer, and can be downloaded into laptop via Consensys software [^3] https://shimmersensing.com/product/consensyspro-software/ for further data processing. A camera was used to record the experiment for annotation. The sensor deployment and data collection example are shown in Fig. 1.
@@ 3.2 Full-Day Data Collection
This research was approved by the ethical committee of KU Leuven with project number: G-2021-4025-R4. Informed consent was obtained from all participants. Two datasets were collected for this study, namely the well-annotated full-day-I (FD-I) dataset, and holdout full-day-II (FD-II) dataset. There is no participants overlap among the two datasets. The statistics of the datasets are shown in Table 1.
@@ 3.2.1 FD-I Dataset
The FD-I dataset contains 34 days of IMU data from 34 participants (6 of them are from our previous study on drinking activity detection [34]). On the data collection day, our research assistants met the participant, instructed the participant to wear IMU wristbands. Participants were free to engage in their normal daily activities. Our research assistants were responsible for the recording when participants change their locations to ensure that all eating and drinking gestures were captured. Specifically, participants were accompanied by one of our research assistants. To clarify the data collection protocol, we have formulated the participants’ daily activities into two states, as illustrated in Fig. 2(a). In state 1, the participant was engaged in activities (e.g., studying, working, and taking meals) at a specific location (Location A). In state 2, the participant was transiting from Location A to Location B by walking, biking, or taking a bus. During state 1, the camera recorded continuously to ensure all intake gestures were captured, especially outside of meal sessions. However, during state 2, we did not record under the assumption that no intake gestures occur during the transition. Specifically, recording started upon the participants’ arrival at Location A and stopped upon their departure. No recording took place during their transition from Location A to Location B. A new recording started upon their arrival at Location B. The assistant mainly used a Logitech camera connected to a laptop to record video during state 1. There were no restrictions placed on participants’ activities or locations during data collection. Using a small photography tripod, the orientation and height of the camera could be easily adjusted to facilitate the deployment. At the beginning of every video recording, the timestamp was presented to enable the synchronization of IMU and video data. At least one meal was collected from each participant, and the minimum data collection duration was 6 h. Both solitary and social eating scenarios were included in the dataset. Participants received a restaurant voucher (20 euro) as experiment compensation after the data collection. A total of 251.70 h two-hand IMU data were collected, which contains 74 eating episodes, with 4,568 eating and 1,100 drinking gestures. The dataset contains four eating styles including fork & knife, chopsticks, spoon, and hand, as shown in Fig. 3(a). Eating locations in this dataset include participant’s home (apartment, student residence), restaurant, and workspace (library, university learning center, and campus rest areas), as shown in Fig. 3(b).
According to the data collection protocol, we recognize that the experimental conditions differ from controlled meal sessions, which are characterized by a single location and short duration. Our approach involves extended monitoring periods and multiple locations within participants’ daily routines. However, these conditions do not fully represent free-living environments as a research assistant consistently accompanies the participants during data collection to ensure the capture of all intake gestures. This presence may introduce additional interference in participants’ daily activities. Therefore, we employ the term near-free-living environments to describe our experimental conditions.